### v2

# Automatic Intermodal Loading Unit Identification using Computer Vision: A Scoping Review

Abstract
Background: The standardisation of Intermodal Loading Units (ILUs), including containers, semi-trailers, and swap bodies, has transformed global trade, yet efficient and robust identification remains an operational bottleneck in ports and terminals.
Objective: To map Computer Vision (CV) methods for ILU identification, clarify terminology, summarise the evolution of proposed approaches, and highlight research gaps, future directions and their potential effects on terminal operations.
Methods: Following PRISMA-ScR, we searched Google Scholar and dblp for English-language studies with quantitative results. After dual reviewer screening, the studies were charted across methods, datasets, and evaluation metrics.
Results: 63 empirical studies on CV-based solutions for the ILU identification task, published between 1990 and 2025 were reviewed. Methodological evolution of ILU identification solutions, datasets, evaluation of the proposed methods and future research directions are summarised. A shift from static (e.g. OCR-gates) to vehicle mounted camera setups, which enables precise monitoring is observed. The reported results for end-to-end accuracy range from 5 % to 96 %.
Conclusions: We propose standardised terminology, advocate for open-access datasets, codebases and model weights to enable fair evaluation and define future work directions. The shift from static to dynamic camera settings introduces new challenges that have transformative potential for transportation and logistics. However, the lack of public benchmark datasets, open-access code, and standardised terminology hinders the advancements in this field. As for the future work, we suggest addressing the new challenges emerged from vehicle mounted cameras, exploring synthetic data generation, refining the multi-stage methods into unified end-to-end models to reduce complexity, and focusing on contextless text recognition. Addressing these challenges can result in minimised human errors, decreased turnaround times, reduced operational costs and higher resilience against disruptions.
keywords:
intermodal loading unit identification , container code recognition ,
journal: Transportation Research Part: C
ILU
Intermodal Loading Unit
TU
transportation unit
UIC
International Union of Railways
UAV
unmanned aerial vehicle
CER
character error rate
IoU
intersection over union
mAP
mean average precision
AP
average precision
FPS
frames per second
GFLOP
giga-floating point operations per second
TRUDI
TRansportation Unit Detection and Identification
PRISMA-ScR
Preferred Reporting Items for Systematic Reviews and Meta-Analyses extension for Scoping Reviews
PCC
Population–Concept–Context
DIP
Digital Image Processing
ML
Machine Learning
DL
Deep Learning
CV
Computer Vision
OCR
Optical Character Recognition
ACCR
Automatic Container Code Recognition
SVM
Support Vector Machine
MLP
Multi-Layer Perceptron
CNN
Convolutional Neural Network
GAN
Generative Adversarial Network
SIFT
Scale-Invariant Feature Transform
HoG
Histogram of Gradients
[…]
[cv]organization=University of Hamburg,addressline=Department of Informatics, Computer Vision Group, city=Hamburg, country=Germany
[…]
[dos]organization=University of Hamburg,addressline=Department of Informatics, Distributed Operating Systems Group, city=Hamburg, country=Germany
[…]
[tuhh]organization=Hamburg University of Technology,addressline=Institute of Maritime Logistics, city=Hamburg, country=Germany
1 Introduction
The term Intermodal Loading Unit (ILU) has been introduced to describe units designed to be easily transferred between different modes of transport, such as vessels, trains, and trucks, without the need to unload and reload the contents. The loading units, such as containers, interchangeable bodies (swap bodies) and cranable semi-trailers, fall under this term [1]. A unique identification code, following the ISO6346 [2] standard, is assigned to each ILU for tracking purposes. This code consists of four letters, a six-digit number and a check digit, which together indicate the owner, type and serial number of ILU. Identically looking ILUs can exclusively be distinguished by their ID codes.
While the standardised identification of ILUs provides a framework for tracking, the scale of modern port operations introduces challenges. As global trade expands, many leading ports handle tens of millions of containers annually, with some operating at or above their nominal design capacity [3]. In this high throughput, keeping track of ILUs is important to ensure efficient management [4]. For example, at transfer points of the terminal to other modes of transport, each ILU must be identified and cross-checked with its associated documentation before entry or exit. When automated systems fail to achieve reliable ILU identification, manual verification becomes necessary, further slowing down the process. Even a brief slowdown at a busy gate can cause truck queues stretching hundreds of metres, disrupting yard operations and increasing turnaround times for transport vehicles [5, 6]. Extended waiting times contribute to congestion which increases fuel consumption and greenhouse gas emissions from idling trucks, amplifying the environmental footprint of port activities [7]. Therefore, efficient and automatic ILU identification is essential for overcoming these bottlenecks and sustaining the growth of containerised transport.
Standardisation and digitalisation enabled the development of various techniques for ILU identification. Among these, Computer Vision (CV)-based identification code recognition stands out as a cost-effective alternative to RFID, which achieves only about 70 % accuracy and entails high installation and maintenance costs [8, 9]. In contrast, CV-based methods can be implemented using a range of camera setups as shown in Figure 1. From static Optical Character Recognition (OCR) gates [10, 11], that can identify entering and exiting ILUs [12, 13], the field is shifting towards more flexible options such as camera equipped aerial or ground vehicles, or even mobile devices, which significantly reduce the need for specialised infrastructure [14, 15, 16, 17]. Furthermore, CV-based ILU identification plays a crucial role in enhancing security by providing accurate verification for customs clearance, which can support in the correct declaration of goods [18, 19, 20].
Figure 1: Schematic overview of a terminal, highlighting the use of camera-equipped aerial and ground vehicles and OCR-gates for comprehensive port monitoring.
Despite its crucial role in improving port operations by reducing human-related errors and increase the processing speed [21, 22, 23, 24], to the best of our knowledge, there are not any literature reviews on CV-based ILU identification. To fill this gap, we prepared this literature review by following Preferred Reporting Items for Systematic Reviews and Meta-Analyses extension for Scoping Reviews (PRISMA-ScR) guidelines [25] with the intention of identifying the available publications and proposed concepts.
In summary, there are 4 contributions of this review:
  • 1.
    Catalogue all empirical studies on ILU identification using CV methods between 1990 and 2025.
  • 2.
    Re-evaluate the terminology used in ILU identification field and provide clear definitions.
  • 3.
    Synthesise an overview of ILU identification methods, image acquisition settings, datasets, and evaluation metrics used in these studies.
  • 4.
    Highlight research gaps and suggest future priorities.
2 Terminology
As CV-based ILU identification is an interdisciplinary field we would like to establish a common terminology to create a common ground for researchers with different background, such as robotics, artificial intelligence, and logistics. In the related literature, we realised a lack of definitions which results in ambiguity. For example, text spotting, an integrated task encompassing both text detection and text recognition, is commonly reduced to OCR, ignoring its broader scope. This ambiguous terminology complicates result interpretation, and hindering meaningful cross-study comparisons. Therefore, we provide the precise definitions for general approaches and downstream tasks.
2.1 Approaches
Digital Image Processing (DIP): Classical, rule-based image operations (e.g., denoising [26], contrast normalisation [27], edge detection [28, 29], morphologic operations [30], thresholding [31, 32], geometric transforms [33]). These operations can be further extended for extracting hand-crafted features, such as edge detection [34, 28], Scale-Invariant Feature Transform (SIFT) [35] or Histogram of Gradients (HoG) [36]. It requires no training but is sensitive to illumination, motion blur, occlusions, or viewpoint changes [37, 38, 39]. The use cases of DIP methods in ILU identification are image enhancement [40], text localisation and character segmentation [41, 42], without relying on data-driven patterns.
Machine Learning (ML): While DIP methods rely on fixed, hand-crafted rules and deterministic operators (e.g. edge detectors, thresholding, morphologic filtering), ML introduces adaptability by learning from data. ML methods use features engineered via DIP (such as HoG, SIFT) as input to classifiers that can discriminate among classes (e.g. character classes, object vs. non-object, scene categories). Common models include Support Vector Machines [43], Random Forests [44], decision trees [45], k-Nearest Neighbors [46] and Reinforcement Learning [47, 48, 49]. These reduce manual rule-tuning and can generalise better under variations in viewpoint, scale, and lighting, but still depend critically on quality of the features and consistency in preprocessing.
Deep Learning (DL): A subfield of ML using deep neural networks, such as Convolutional Neural Networks [50, 51] or Transformers [52] that learn features and classifiers jointly and end-to-end. Although DL requires substantial training data and computational resources, it automates the feature engineering process. However, a major drawback of deep learning models is their black-box nature, which makes them difficult to interpret and explain [53]. For similar tasks depending on similar features, deep learning methods can benefit from transfer learning [54] and fine-tuning [55].
2.2 ILU Identification Related Computer Vision Tasks
Besides the terminology for general approaches, we would like to provide definitions for downstream tasks:
Object detection: Localising (drawing bounding boxes around) and classifying (assigning labels to) objects within an image. Unlike image classification (which labels are for the entire image), object detection handles multiple objects per image, providing both what and where the objects are [56, 57].
Instance segmentation: Localising (pixel-level regions, drawing polygons around) and classifying individual objects within an image. Unlike object detection, which uses bounding boxes, instance segmentation assigns pixel-level class labels and provides precise shapes of the objects [58].
Character segmentation: Localising and segmenting characters in an image, typically as a prior step for recognition [59].
Character recognition: Classifying a single already localised character into its alphanumeric label. This problem is often tackled as an image classification task where the input is a cropped character patch and the output is the label for this single character patch [60].
Character detection: Localising (drawing bounding boxes around) and classifying (assigning labels to) characters within an image. Similar to the object detection, character detection provides information about where individual characters are and which characters they are [61].
Optical Character Recognition (OCR): Localising and identifying individual characters in an image. OCR is predominantly used for document text recognition and perform well under standard conditions, such as uniform backgrounds and consistent lighting (e.g., scanned documents). However, OCR techniques may fall short when dealing with text-in-wild scenarios [62], where background and lighting are highly diverse and text orientation is arbitrary. Relying on character detection in diverse scenes can increase false positives, as many objects share similar patterns with individual letters [63, 64]. Another disadvantage of character-level recognition, especially with DL methods, is that character-level recognition needs more expensive annotated data than word-level or line-level recognition due to its granularity and required labour for annotating individual characters.
For tasks involving the diverse scenes, such as identifying signs in urban areas [65], two primary subtasks are defined:
(Scene-)Text detection: Localising regions containing text in an image without extracting its textual content. Outputs quadrilaterals around words or lines. Unlike character detection, it focuses on full words or lines, which makes it more robust against distractors [66, 67, 68, 69].
(Scene-)Text recognition: Extracting textual information from detected text regions by converting them into machine-readable characters or words [70, 71, 67, 68, 69, 69].
Text spotting: End-to-end method that jointly detects and recognises text in a single step [72, 61].
2.3 Terminology Beyond OCR Gates
Terminal entrances remained as the critical bottlenecks in terms of operational efficiency and congestion impact, so their optimisation becomes one of the key challenges in terminal management. The earlier studies on this, applied CV methods only at terminal gates, where the task is to automatically identify loading units as they entered or exited the facility. In the literature, this task is referred to as Automatic Container Code Recognition (ACCR) [73, 74, 75, 76, 77, 78, 79] and rarely as loading unit identification or transportation unit (TU) identification [17]. The focus of ACCR is on containers at terminal gates. However, the word container does not cover all types of loading units such as semi-trailers and swap bodies that has ISO6346-compliant codes (as defined in EN 13044 [80]). Second, the word recognition does not fully capture the actual tasks involved.
The term recognition refers to extraction of text from an image in the CV field. However, this presupposes that the visual input image consists solely of the target text. This assumption is rarely valid in the real-world scenarios involving surveillance cameras or mobile cameras, where text is typically embedded within complex and cluttered scenes. In these cases, an additional stage, text detection is required to localise the text regions before recognition can be performed [81]. Over time, these methods were extended beyond the gates. With advancements in CV and robotics, identification is now performed at quay cranes [82], yard blocks, and terminal vehicles such as reach stackers [83]. These new use cases introduce more variation and complexity than the original gate-based setting and therefore, text recognition is not sufficient by itself.
For these reasons, we propose the term Intermodal Loading Unit Identification as an inclusive designation for the task of determining ID codes of ILUs. This term covers various loading units and also ILU detection/segmentation stage prior to text detection and recognition steps, introduced in some studies [17, 84]. A term that covers all these stages aligned with CV terminology is important, as it would address shared challenges across the CV applications. Using a common terminology can link the relevant studies to the broader problems. For example, different scales of objects and text is a problem that was introduced with the mobile cameras for ILU identification task. However, small object segmentation [85] or small text detection [86] havebeen actively studied in general purpose CV methods. Moreover, a common terminology can help in finding the distinct aspects of ILU identification from the broader tasks. Unlike in common text recognition, where a language model based encoder (e.g. TrOCR [87]) is useful due to contextual cues present in the natural language, the contextless nature of ISO6346-compliant ID codes makes this approach ineffective. Such models tend to overfit to frequently occurring text instances in training data [15], limiting their applicability in this context. Adopting the term ILU identification can provide conceptual clarity and establish a coherent basis for researchers and practitioners.
Having established the terminology and challenges, we now turn to the methodology used for systematically review the literature.
3 Methodology
This literature review was conducted in accordance with the PRISMA-ScR guidelines [25] and eligibility criteria were specified a priori and piloted before the screening began. These criteria follow the Population–Concept–Context (PCC) logic and Table 1 details the inclusion and exclusion rules together with their rationales.
Table 1: Inclusion and exclusion criteria for systematic review of ILU code recognition.
Domain Inclusion rule Exclusion rule
Date range 1 Jan 1990–31 August 2025 << 1990 or after cut-off
Language English full text Non-English; abstract-only
Publication type Peer-reviewed journal / conference paper; pre-print (quantitative) Patents, standard, theses, news or editorials
Population (P) ILUs (containers, swap-bodies, semi-trailers) Other objects unless ILUs also analysed
Concept (C) CV localisation and recognition of alphanumeric ILU code RFID/barcode solutions
Context (C) Fixed or mobile cameras in ports, terminals, rail yards Depots
Reporting completeness Method detail plus numeric outcomes (precision, recall, F1, mAP, read-rate) No quantitative results
Duplicates / accessibility First complete, accessible full text Duplicates; inaccessible full text
As the first review on ILU identification, we applied inclusive eligibility criteria to capture the full scope of available evidence. Our review includes articles written in English, published from 1990 to 2025, that follow a scientific structure. Both peer-reviewed articles and preprints are included as long as they report quantitative results, allowing us to catalogue the latest research findings. The final selection of articles and data charting were determined through consensus among the authors, following a review of titles, abstracts, and full texts.
We finalised our search for articles in August 2025, using Google Scholar and dblp-Computer Science Bibliography to find scientific articles on ILU identification using CV or DIP techniques. Our search terms included container code recognition, trailer code recognition, and identification of containers and trailers. As discussed in Section 2, the terminology used in this field is not standardised and it directly affects our search methodology. We checked the cited papers and ’related articles’ in Google Scholar and Research Rabbit to expand our pool of eligible articles.
From the 63 eligible articles the following data items were extracted into a spreadsheet [^1] Available in supplementary materials.: journal or conference details, contributions, methods used, method class (DIP, ML, or DL), dataset specifics (availability, number of images, diversity in lighting and weather), image collection region and camera position such as fixed (CCTV) or mobile (aerial or ground vehicle), evaluation metrics, results, deployment status, suggested future directions and funding. The resulting table was then analysed and discussed by the authors.
4 Results and Discussion
Our search yielded 91 articles published between 1990 and 2025. During the screening process, articles were excluded due to being duplicates ($n=2$), not reporting quantitative results ($n=6$), written in a non-English language ($n=5$), focusing solely on container detection or integrity assessment rather than identification via ISO6346 compliant codes ($n=6$), solutions for other industrial text recognition applications such as wagon code recognition ($n=5$), unavailability of their full text ($n=2$), or being patents ($n=2$).
Figure 2: Geographic distribution of articles by country (n=63). Five articles involve international collaborations, which results in 69 for country mentions rather than 63.
Most of the published articles were written by the authors with Asian affiliations (79,71 %) as shown in Figure 2. This is related with the high volumes of ILU handling in Asian ports and terminals, which creates a need for efficient ILU identification systems. Countries such as China, Taiwan, South Korea, Vietnam, and Japan play significant roles in global trade. This leads to innovation as reflected by the contribution of Asian countries to the scientific publications. Substantial investment in research and development by Asian governments and companies further supports this innovation [88, 89].
Public funding is reported as the main source, supporting about two out of five publications, indicating strong governmental backing for the logistics sector that operates with low profit margins [90, 91]. The motivation for strong governmental backing is the potential for technological advancements that will benefit national infrastructure and economic growth [92, 93]. Besides this dominant public sector support, industry funding and public-industry collaborations constitute one ninth of the publications, suggesting that while there is some industry involvement, companies largely rely on public sector investments to drive technological development of ILU identification systems. Nevertheless, the presence of pure industry funding (4,76 %) indicates a genuine market demand for companies to invest in research to remain competitive [94, 4].
Figure 3: Ranking of publication outlets for the eligible articles ($n=63$) over the year. The years with zero eligible publications are ommited in this barchart. As this scoping review was prepared in 2025, the data for that year remain incomplete, covering only publications up to August.
While funding sources indicate industry and governmental priorities, the ranking of publication venues offers insight into the field’s academic recognition. To analyse the rankings of journals and conferences we followed the rankings of Scimago [^2] https://www.scimagojr.com/ for journals and CORE [^3] https://www.core.edu.au/icore-portal for conferences. Figure 3 shows that more than half of the articles on automatic ILU identification task are published in unranked outlets, which is not uncommon for applied or niche technical topics [95]. Especially, the early years of DL dominance marked a period of stagnation (2017-2019), where the field struggled to deliver methodological innovation. This trend is changing in the recent years as the percentage of articles published in ranked outlets are increasing, as well as the total number of publications, suggesting that the ILU identification topic is relevant for high-quality research venues. Nevertheless, the field could benefit from a greater emphasis on more novel contributions, increased comparability, and reproducibility of the results. This could be achieved through publicly available benchmark datasets and repositories that showcase the source code and the model weights. By adopting such research standards, ILU identification tasks with CV-based methods could be more relevant for high-ranked publication outlets.
4.1 Methods
The development of methods for ILU identification has undergone a significant transformation over three decades, as illustrated in Figure 4. For two decades between 1990 and 2010, the field relied predominantly on DIP techniques combined with early ML methods and to a lesser degree, DIP methods alone. This mirrors broader trends in the CV literature [96]. In this period, researchers mainly followed a three-stage approach: 1) Removal of the background using thresholding [97, 41], edge detection [41, 98] and morphological operations [99, 98]. 2) After segmenting the characters, each was classified by methods like template matching [100, 101, 102, 98], or traditional ML methods like Multi-Layer Perceptron (MLP) [103, 99, 42] or fuzzy logic [104]. 3) Finally, the predictions on the individual characters were postprocessed to output a full ID code. These methods performed adequately in controlled image acquisition systems like OCR-gates [105] but struggled with challenging conditions, such as variability of brightness [75] and damages [106, 107] on the ILU images.
Figure 4: The number of publications and their approach categories used for ILU identification task over the years.
In the subsequent five years, a modest increase in ML-based methods can be observed, particularly in hybrid DIP and ML approaches [108, 109, 110, 111, 112, 8]. This reflects the growing recognition of ML’s potential for enhancing robustness in character and symbol recognition tasks in CV [113]. However, the field lagged several years behind the broad adoption of DL, despite its transformative influence on CV solutions [114, 115] from 2012 onwards (e.g., the emergence of AlexNet [116]). This delay may be attributed to the scarcity of labelled ILU identification datasets, limited computational infrastructure, lack of technology skills in the workforce, and the conservative approach for adopting new technologies in the logistics industry [117, 118].
The five-year interval from 2016 to 2020 marks a turning point, with a notable adaptation of DL-based methods. Publications using only DL-based approaches rise sharply, as well as the hybrid strategies combining DIP, ML, and DL. This phase was driven by advances in GPU hardware [119, 120], and the availability of open-source frameworks such as TensorFlow [121] and PyTorch [122]. Deep architectures, particularly CNNs [50], rapidly proved their superiority for conducting ILU identification under challenging conditions [79, 123]. As a consequence, purely DIP-based approaches declined significantly, often limited to pre- or post-processing roles within hybrid pipelines [75, 84, 21, 124, 77].
Table 2: Stages defined in DL-based pipelines for ILU identification. The table reports how often each combination appears in the surveyed literature. $\bullet$ = detection, $\bullet$ = recognition.
Object Character Text Count Percentage
20 54,1%
4 10,8%
3 8,1%
2 5,4%
2 5,4%
2 5,4%
1 2,7%
1 2,7%
1 2,7%
1 2,7%
Until 2016, all methods developed for ILU identification focused on character-level localisation and recognition with fixed camera settings. This approach is still in use with DL methods [16, 125, 126] as ILU identification on fixed cameras often resembles OCR on scanned documents. However, with the increasing use of vehicle-mounted cameras, the task has evolved to resemble scene-text detection and recognition rather than OCR. As a result of this, the trend is shifting towards using comprehensive scene-text detection and recognition methods rather than relying on character detection and recognition. Table 2 shows that the majority of articles employ DL methods for both scene-text detection and recognition tasks. The share of hybrid approaches using non-DL methods for one of these steps are minimal. Examples include using traditional techniques for text detection such as cropping a predefined area for searching ID codes [127] or for character recognition, template matching after a DL model detects text areas [128]. Instead, researchers largely favour more comprehensive pipelines that rely on DL in each step.
By 2021–2025, DL-based approaches had established themselves as the predominant paradigm for ILU identification. Even hybrid DIP-based methods virtually disappeared, indicating that the DL models now offer both the required accuracy and operational robustness [73, 129, 130]. The decline of hybrid approaches suggests that most effective solutions now rarely require DIP or traditional ML methods, except in specialised scenarios such as real-time constraints [106, 13, 19, 127, 126, 12]. Researchers also address real-time constraints by focusing on text spotting as these methods are generally faster and more efficient due to their unified feature extraction module [131, 132, 133]. A few studies introduce object detection or instance segmentation as a prior step to text detection and recognition [17, 84]. Detecting the ILU reduces the search area for the text detection step and enables association of detected and recognised IDs with other objects, thereby improving scene understanding [17].
4.2 Datasets
The quality and diversity of annotated data strongly influence how well DL models perform once deployed in real-world port environments. Carefully curated datasets with consistent labelling not only enable robust training and evaluation, but also determine whether methods can generalise beyond laboratory settings. Understanding the properties of datasets is therefore essential for evaluating the performance and applicability of CV-based ILU identification systems. Below, we summarise the main characteristics of the datasets used in the eligible studies, highlighting camera position settings, diversity, and availability.
4.2.1 Image Acquisition Settings
We categorised cameras as fixed or mobile cameras based on their mounting and usage. Fixed cameras include OCR gates, CCTV cameras, and crane-mounted cameras. These are installed at permanent locations such as terminal gates, yards, or crane structures, providing stable, controlled views for capturing ILU codes. Their main advantage is delivering consistent, high-quality images ideal for computationally efficient methods like classical DIP and traditional ML, as long as ILUs are present at predefined areas. Although the installation and operation of these systems involve substantial infrastructure investment, fixed cameras often struggle to capture sufficient feature information in large and complex environments, resulting in reduced accuracy and reliability of detection and recognition based on visual neural networks [134].
Figure 5: Sample images for fixed perspective (left) [79], ground vehicle perspective (middle) [17], aerial perspective (right) [15].
In contrast, mobile cameras are mounted on unmanned aerial vehicles, reach stackers, terminal trucks, or other industrial vehicles and smart devices. For ground-level perspectives, we also considered handheld cameras, as they closely resemble those captured by vehicle-mounted systems. Unlike fixed setups, mobile cameras provide flexibility by moving across the terminal and capturing a wider range of perspectives. This mobility introduce higher image variability, requiring advanced processing methods and diverse datasets. When these challenges addressed, there is a potential for more precise and real-time terminal monitoring. Thanks to sensor equipped vehicles [135, 83], CV pipelines beyond ILU identification can be created. For example, localisation of ILUs through pose estimation, autonomous path optimisation for vehicles navigating around terminals, multi-agent coordination (collaboration of aerial and ground vehicles) for terminal monitoring.
Table 3: Summary of dataset characteristics: camera position and dataset availability.
Category Count Percentage
Camera position Only Fixed 47 74,60 %
Fixed and Ground 10 15,87 %
Fixed, Aerial and Ground 3 4,76 %
Only Ground 1 1,58 %
Only Aerial 1 1,58 %
Not reported 1 1,58 %
Availability Not public 54 85,71 %
Fully available 3 4,76 %
Upon request 3 4,76 %
Partially available 2 3,17 %
Behind paywall 1 1,58 %
As shown in Table 3 the most common collection setting is the fixed CCTV cameras at OCR-gates [136, 125, 137] or cranes (74,60 %) [84, 110] and a mixture of these images with smartphone images (15,87 %) [74]. However, vehicle-mounted cameras are increasingly used alongside fixed cameras to capture wider areas and provide information across the terminal [14, 15, 16, 17].
4.2.2 Diversity
Most datasets consist of images gathered from ports, reflecting their operational settings. Images in these datasets were taken from the world’s busiest ports like Shanghai [13], Singapore [42], Hong Kong [138], Kaohsiung [9], and Hamburg [17], as well as other ports in Malaysia [73], Sweden [125], Vietnam [19, 126, 127, 75], Thailand [21], South Korea [14, 78, 139, 110, 104], Japan [101], and Italy [140]. While rare, there are datasets with images collected in smaller inland ports like Dresden and Riesa [15, 17]. Besides the terminal-specific data collection, web scraping were also used [129, 79]. In addition, controlled environments such as the Shanghai Maritime University test facility [40] were used for data collection. However, datasets consisting of images from multiple locations are rare. For DL methods, capturing the nuances of ILUs across different countries and their respective ports is important.
The number of images used in the experiments in the eligible articles ranges from a minimum of 30 images to a maximum of 34.000, with a median of 1.050 images. Since the DIP-based methods do not require a large amount of data, researchers who tackled ILU identification task before DL, experimented with a lower number of images compared to recent studies. However, with the emergence of ML and DL, the need for data has increased significantly. Most of the studies have addressed this need by pre-training their models on general-purpose scene-text detection and recognition datasets such as ICDAR2015 [141], CTW1500 [142], and SynthText [143], then fine-tuning on smaller ILU datasets [16, 17].
Starting from the early datasets, efforts to represent diverse conditions have been made in order to evaluate generalisation. Lighting variability is the most discussed factor, with datasets often including images taken in bright, low-light, and uneven lighting scenarios [102, 41]. Considering how the earlier, thresholding-based methods get affected by small changes in light intensity, diversifying the lighting conditions in the experiments was a major focus [140, 99, 138, 98]. However, currently lighting conditions are classified with more general classes: Day and night settings are frequently included to test the adaptability of the systems [106, 16, 17, 13].
Weather conditions such as rain, snow, and fog are also represented, as they can significantly affect image quality and ILU identification performance [17, 74, 130, 138]. Additionally, some datasets feature images with motion blur, camera noise, dirty, rusty, damaged, or partially occluded ID codes to further challenge algorithms with the conditions that exist in real-world settings [106, 17, 107, 130, 79]. Rarely were these conditions simulated to increase the diversity of datasets [129].
4.2.3 Availability
We divided datasets into 5 different categories: Fully available [24, 107, 17], upon request [144, 16, 136], partially available [145, 79], behind paywall [14] and not public. Table 3 shows the availability of the datasets used by the identified articles. The significant majority of the datasets are not publicly available, which creates a substantial barrier for research and development in this area. We were not able to confirm the availability of the datasets that claim to be available upon request as our requests went unanswered. Recent papers acknowledge the lack of benchmark datasets and its negative effects on the field [16, 17]. The lack of availability prevents the reproducibility of results and fair comparison between the proposed methods for researchers. As for the practitioners, while large seaports typically have extensive propriety datasets, publicly available datasets extend these capabilities to smaller terminals, facilitating data-driven optimisation.
After screening the literature, we could find only 3 publicly available datasets. Among these, the TRansportation Unit Detection and Identification (TRUDI) dataset [17] stands out for its annotation density. It offers 35,034 instances across 733 images, covering five classes: Container, trailer, tank container, ID code, and logo. It introduces motion blur as an additional challenging condition, as most of its images were captured via UAVs, reach stackers, and terminal trucks. The Shipping Container Code dataset [24] leads in the number of images, and the Container Number-OCR dataset [107], though smaller (2.851 images from Tianjin Port), offers similar challenging conditions as others: variable weather conditions and degraded ID codes (dirt, rust, damage). Further details for the publicly available datasets are shown in Table 4.
Adopting these publicly available datasets for benchmarking will be very helpful for both researchers and practitioners to evaluate and compare proposed methods.
Table 4: Publicly available ILU identification datasets.
Dataset Images Instances Conditions Location
Shipping container code [24]
3600
4000
Day/night, Rainy, snowy Dirty, damaged ID Not reported
Container Number-OCR [107]
2851
2851
Low/high light Foggy Dirty, damaged ID Tianjin Port
TRUDI [17]
733
35034
Day/night Rainy, snowy, foggy Motion blur Dirty, damaged ID
Multiple countries,
Multiple terminals
4.3 Evaluation
To evaluate the proposed ILU identification methods, several evaluation metrics are commonly used. As there are multiple stages of ILU identification using CV methods, such as text detection and text recognition, each stage can be evaluated individually. In the literature, 60.18 % articles reported the results of individual stages only and did not include end-to-end results. However, the results for individual stages offer limited insight for the applicability of the method for the ILU identification task. End-to-end accuracy or success rate is the most relevant metric due to its practical importance. It is the most strict metric as one misrecognised character is enough for failed identification.
Precision and recall are the common metrics for these individual stages. These metrics show the balance between false positives and false negatives, with precision focusing on the accuracy of positive predictions and recall on the model’s ability to identify all relevant instances. The F1-score, or H-mean, harmoniously combines precision and recall into a single metric, offering a balanced view of the model’s performance.
For object detection, instance segmentation or text detection tasks, average precision (AP) and mean average precision (mAP) are commonly used, measuring precision across various intersection over union (IoU) levels. The IoU metric is important for assessing the overlap between predicted and ground truth bounding boxes. As for the recognition, besides precision, recall and F1-score, other metrics such as the character error rate (CER) or word accuracy are used. CER measures the proportion of characters that were incorrectly predicted compared to the total number of characters in the reference text, using the Levenshtein-distance [146].
Speed and efficiency are evaluated using metrics such as frames per second (FPS), and inference time, which assess the model’s operational speed. Model complexity is assessed through the number of parameters, giga-floating point operations per seconds, and model size, indicating the computational demands and storage requirements.
Although the recent studies use the same, established metrics adapted from general purpose text detection and recognition task, putting the reported results for ILU identification into a frame is not possible with the lack of an widely adopted benchmark dataset and the lack of end-to-end results. Only around 40 % of the studies reported end-to-end accuracy and it ranges from 5 % [14] to 99,51 % [74], with the mean of 84.19 % and standard deviation of 24.02 percentage points. Considering the high variation and the low number of reported end-to-end results ($n=25$), we did not perform further statistical analysis as these will not yield in statistically significant results.
4.4 Future Research Directions
In the literature, future directions for the ILU identification task mainly focus on enhancing the datasets by including more images taken from diverse environments, introducing different challenges such as various distances, angles, weather and lighting conditions. This expansion aims to cover multiple ports and scenarios to improve model generalisation. However, increasing the diversity of proprietary datasets would not solve the fundamental problem with the datasets for ILU identification. Given that creating a new dataset for downstream CV tasks are tedious, time consuming, and costly [147, 148], alternatives such as synthetic data generation can be considered [149, 143], e.g., by using a game engine [150].
Handling challenging conditions, such as low-light scenarios, occlusions, and degraded or damaged ID codes, is another direction pointed by the researchers. Suggested methods include image enhancement techniques such as noise reduction [18], or the use of Generative Adversarial Networks [151, 152] to address issues such as shadows, rust, and dirt [24, 107]. Super-resolution [153] is a task that can tackle the problems with degraded or damaged ID codes. However, the literature on scene-text super-resolution has focused primarily on natural language, leaving an opportunity for research specifically on contextless text such as ILU codes. Alternatively, multi-view ILU identification is another future direction due to its potential against degraded ID codes [24, 8] as multiple images of a single ILU are captured from different angles in this approach. Postprocessing steps are also recommended to verify recognition results by using the check digit and the owner code. [12, 154, 76, 79].
For DL-based methods, refining network architectures to increase the receptive field so that the model can see more of the scene, especially smaller objects in a single pass [16]. Another future work suggestion is incorporating attention mechanisms to boost recognition accuracy [155] or optimising the tokenisation process of vision transformer-based architectures as there are always the same amount of characters expected for ISO6346-compliant ID codes. For better applicability, the development of unified, end-to-end models is suggested to streamline processing pipelines [17, 136, 107, 154]. While DL models achieve high accuracy, their computational demands often limit deployment on edge devices like drones. Future work should focus on model compression, quantisation, or neural architecture search to enable real-time ILU identification in resource constrained environments.
The contextless nature of ISO6346 codes makes scene-text spotting task harder as there are no natural language cues [156]. Therefore, directly applying current scene-text recognition models that rely on semantic context, is not the most suitable approach, as they usually use a language model or at least a decoder module pretrained on natural language to correct the initial recognition result [157, 87]. However, incorporating complex language models increases model complexity, while not yielding any benefits for contextless recognition tasks [156]. Researching on the contextless scene-text recognition models is an opportunity which may result in novel and more efficient methodologies. These new methods can also be useful for similar applications including International Union of Railways (UIC)-wagon codes, or registration plates.
5 Conclusion
This review provides the first systematic synthesis of CV-based ILU identification research by analysing 63 empirical studies published over three decades, and propose a more inclusive terminology. There is a clear upward trend for the quality and number of publications in the recent years driven by the adaptation of CV-based systems, which reduce human error and increase efficiency.
Over the years, the field has shifted from classical DIP to DL thanks to the advancements in computational resources and deep learning frameworks. This shift comes with the need of huge datasets to train DL models and dataset accessibility is the biggest obstacle with 85,71 % of datasets being private. This creates problems for both researchers and practitioners. While large seaports have the means of creating large scale datasets, small sized ports have limitations to create datasets with the same scale, which creates a competitive disadvantage. The median dataset size is too small for robust DL training, forcing researchers to adapt general purpose text detection and recognition models, causing text recognition models to overfit on words with semantic meaning. Limited dataset accessibility also affects the quality and credibility of publications as there are no adapted benchmark datasets. This makes performance comparisons difficult along with the lack of end-to-end analysis, leaving gaps in understanding which methods perform best under which setting. To the best of our knowledge, there are only three publicly available datasets which can be adopted for benchmarking. Considering that two of them were published within 2025, in the near future, the adaptation of these datasets might resolve this problem.
While the rise of DL-based methods brought the need for publicly available datasets, but the shift from static to dynamic image acquisition settings brings other dimensions. The research on ILU identification relied on fixed camera settings such as OCR-gates. However, recent studies are increasingly adopting vehicle mounted camera settings. This shift introduces new challenges like higher variations in object scale caused by changing camera-to-object distances and unknown ILU locations. This is why, other CV and robotics related tasks such as pose estimation (predicting the real-world coordinates and orientation of ILUs) or mission planning for autonomous vehicles (e.g. calculating the most optimal route for identifying ILUs) are now emerging priorities. The seemingly simple problem of identifying ILUs is a fundamental step that has not been taken for autonomous terminal operations. Therefore, addressing these challenges holds transformative potential for transportation and logistics as precise monitoring can optimise the terminal operations, reduce human errors, pave the way for faster turnaround times, lower operational costs and increased resilience against disruptions like misplaces ILUs.